OpenStudy (anonymous):

write the expression as a complex number in standard form (5+4i)+(7+2i)

hartnn (hartnn):

add real parts, 5+7=... ? add imaginary parts 4i+2i =.... ?

OpenStudy (anonymous):

so 13 and 6i?

hartnn (hartnn):

infact 5+7 = 12

hartnn (hartnn):

so, 12+6i